A small helium-neon laser emits red visible light with a power of 5.80 mW in a beam that has a diameter of 2.90 mm .

a) What is the amplitude of the electric field of the light?

b)What is the amplitude of the magnetic field of the light?

c)What is the average energy density associated with the electric field?

d)What is the average energy density associated with the magnetic field?

e)What is the total energy contained in a 1.00-m length of the beam?
Answer #1

a) The intensity of laser beam is (5.8x10-3 w) a(1.45x10-m) = 878.1 W/m2 The amplitude of electric field is 2(878.1 W/m²) (8.

b) The amplitude of magnetic field is Bmas = max C 813.31 V/m 3.0x108 m/s = 2.71x101 (C) & (d) The average every densities as

U Eqv = UBAV = 48.85*10–12 c²N - m?)(813.31 V/m)2 = 1.463x10-6J/m3 The total energy contained in 1.0 m length of the laser is